OpenStudy (anonymous):

Help! Find the exact value of cos(11π/12). Hint: begin by rewriting 11π/12 in terms of familiar angles.

OpenStudy (anonymous):

You are allowed to use calculators right? :/

OpenStudy (anonymous):

Well if you are not, first change 11π/12 into a familiar angle. π=180 degrees in this case because it's a trig type question.

OpenStudy (anonymous):

You can use calculators

OpenStudy (anonymous):

11x180/12 =1980/12 =165 That doesn't seems right.

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Is it correct, because it looks like I made a mistake

OpenStudy (anonymous):

It is correct Red_Stevens

OpenStudy (anonymous):

Really? That isn't any of the answer choices.

OpenStudy (anonymous):

Alright I'm going to resubmit this.

OpenStudy (anonymous):

\[\cos \frac{ 11\pi }{ 12 }=\cos \left( \pi-\frac{ \pi }{ 12 } \right)=-\cos \frac{ \pi }{12}=-\cos 15\] \[\cos 15=\cos \left( 45-30 \right)=\cos 45\cos 30+\sin 45\sin 30=\frac{ 1 }{\sqrt{2} }\frac{\sqrt{3}}{ 2}+\frac{ 1 }{ \sqrt{2} }\frac{1 }{2 }\] \[\frac{ \sqrt{3}+1 }{2\sqrt{ 2} }\times \frac{ \sqrt{2} }{ \sqrt{2} }=\frac{ \sqrt{6}+\sqrt{2} }{4}\] now you can write the solution.